### v1

# Resisting Quantum Key Distribution Attacks Using Quantum Machine Learning

Abstract
The emergence of quantum computing poses significant risks to the security of modern communication networks as it breaks today’s public-key cryptographic algorithms. Quantum Key Distribution (QKD) offers a promising solution by harnessing the principles of quantum mechanics to establish secure keys. However, practical QKD implementations remain vulnerable to hardware imperfections and advanced attacks such as Photon Number Splitting and Trojan-Horse attacks. In this work, we investigate the potential of using quantum machine learning (QML) to detect popular QKD attacks. In particular, we propose a Hybrid Quantum Long Short-Term Memory (QLSTM) model to improve the detection of common QKD attacks. By combining quantum-enhanced learning with classical deep learning, the model captures complex temporal patterns in QKD data, improving detection accuracy. To evaluate the proposed model, we introduce a realistic QKD dataset simulating normal QKD operations along with seven attack scenarios, Intercept-and-Resend, Photon-Number Splitting (PNS), Trojan-Horse attacks Random Number Generator (RNG), Detector Blinding, Wavelength-dependent Trojan Horse, and Combined attacks. The dataset includes quantum security metrics such as Quantum Bit Error Rate (QBER), measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring an accurate representation of real-world conditions. Our results demonstrate promising performance of the quantum machine learning approach compared to traditional classical machine learning models, highlighting the potential of hybrid techniques to enhance the security of future quantum communication networks. The proposed Hybrid QLSTM model achieved an accuracy of 93.7.0% after 50 training epochs, outperforming classical deep learning models such as LSTM, and CNN.
Index Terms:
Quantum Computing, Quantum Key Distribution, QLSTM, QBER, PNS Attack, Quantum Trojan Horse.
Data Availability Statement: Dataset supporting the findings of this study is available from the corresponding author upon reasonable request.
Funding Statement: None.
Conflict of Interest Disclosure: The authors declare that they have no known conflict of interest that could have appeared to influence the work reported in this paper.
I Introduction
The rise of quantum computing poses a significant challenge to the security of modern communication networks, particularly for encryption techniques based on mathematical problems. As quantum computers evolve, the infamous Shor’s algorithm will be able to break today’s asymmetric encryption techniques [1]. In response to these emerging threats, researchers have been working on Post-Quantum Cryptography (PQC) to develop encryption algorithms that can resist quantum attacks. PQC algorithms rely on complex mathematical structures, such as lattice-based, code-based, and multivariate polynomial cryptosystems, making them resistant to current quantum techniques. However, despite its promising security features, PQC remains reliant on mathematical problems, which means that its security is not provably unconditional. Given the rapid advancements in computational capabilities, even these new cryptographic approaches could face vulnerabilities in the future [2].
An alternative and more radical direction is Quantum Key Distribution (QKD), which leverages the principles of quantum mechanics to generate keys that are theoretically secure against eavesdropping [1, 2]. Unlike classical key exchange methods, QKD benefits from the no-cloning theorem, which prevents copying quantum states. This property allows Alice (the sender) and Bob (the receiver) to detect any eavesdropping attempt, as such intrusion inevitably disturbs the transmitted quantum states [3]. This makes QKD attractive for sectors such as government, finance, and critical infrastructure, where extremely high levels of security are required [1].
In recent years, QKD has progressed from theoretical studies to experimental prototypes and early-stage deployments. However, integrating QKD into existing communication infrastructures presents challenges related to scalability, network architecture, and resource optimization [4]. While global-scale QKD networks are being envisioned [5], real-world implementations remain vulnerable to practical threats. Hardware imperfections, channel noise, and side-channel attacks—such as Photon Number Splitting (PNS), Trojan Horse Attacks, and Channel Tampering—can all undermine the theoretical guarantees of QKD [5, 3]. These threats emphasize the importance of robust monitoring and intrusion detection strategies.
Researchers have explored both conventional and data-driven approaches to securing QKD against such vulnerabilities. On the traditional side, non-ML methods focus on signal analysis, hardware countermeasures, and statistical detection. For example, spectral estimation techniques have been used to detect low-rate denial-of-service (LDoS) attacks in CV-QKD by identifying distinct low-frequency patterns in the power spectral density (PSD) [6]. Phase-sensitive amplifiers (PSA) combined with homodyne detection have been deployed to improve resilience against individual attacks [7], while plug-and-play architectures for Differential Phase Shift Measurement-Device-Independent QKD (DPS-MDI-QKD) help mitigate vulnerabilities in state preparation and measurement [8]. Network-layer defense strategies based on real-time traffic monitoring have also been proposed to detect denial-of-service attempts [9]. While these methods enhance QKD robustness, they are often limited by predefined assumptions, making them less adaptable to novel or evolving threats.
This limitation has motivated a growing interest in Machine Learning (ML) and Deep Learning (DL) for QKD security. Unlike static threshold-based methods, ML models can adapt to dynamic patterns and detect both known and previously unseen attacks. In Continuous-Variable QKD (CV-QKD), Du and Huang [10] proposed multi-class detection using Binary Relevance Neural Networks (BR-NN) and Label Powerset Neural Networks (LP-NN), combined with one-class SVM for unknown attacks. Kish et al. [11] developed a decision tree–based framework for detecting and mitigating channel amplification attacks and related DoS scenarios. Mao et al. [12] used Artificial Neural Networks (ANN) to classify multiple attack types from optical parameter features such as quadrature values and local oscillator intensity. Luo et al. [13] introduced a semi-supervised GAN-based method that improved detection generalization to unseen threats. Table I summarizes these CV-QKD approaches.
TABLE I: Detecting quantum attacks in CV-QKD via ML
[13] [12] [10] [11]
Training Data Experimental data from a real CV-QKD system. Simulated CV-QKD parameters. Simulated CV-QKD parameters. Estimated CV-QKD parameters.
Attacks Covered Calibration, LO intensity, saturation, wavelength, unknown threats. Calibration, LO intensity, saturation, hybrid attacks. Calibration, LO intensity, saturation. Channel Amplification (CA), CA-DoS, DoS.
Model GAN with anomaly detection. ANN. BR-NN and LP-NN + one-class SVM. Decision tree.
Accuracy 99.3%. >>99%. 100% (known), 98.7–99.8% (unknown). 100% (low noise), 90.1% (high noise).
Generaliza-tion High Moderate High Low
Advantages Strong generalization. Effective multi-attack classification. Known + unknown detection. Fast classification.
Limitations High complexity. Reduced SKR and distance. Potential false positives. Narrow training scope.
In Discrete-Variable QKD (DV-QKD), similar ML-based methods have been explored. Al et al. [14] applied ANN and LSTM models to detect Man-in-the-Middle attacks in IoT-integrated QKD networks. Tunc et al. [15] used LSTM and SVM classifiers to detect eavesdropping in quantum communication channels. Xu et al. [16] implemented Random Forest classifiers capable of detecting both device imperfections and eavesdropping in real-time. Table II summarizes these DV-QKD approaches.
TABLE II: Detecting quantum attacks in DV-QKD via ML
[14] [15] [16]
Training Data Simulated QKD key length data. Theoretical BB84 attack simulations. Experimental + theoretical data.
Attacks Covered Man-in-the-middle. Eavesdropping. Device imperfections + eavesdropping.
Model ANN, LSTM. SVM, LSTM. RF.
Accuracy 99.1%. 100%. 98%.
Generalization Low Low High
Advantages Effective for predefined attacks. Accurate detection. Detects known + unknown threats in real-time.
Limitations Simple assumptions, high computation time. Limited scope. High complexity.
Despite these advances, existing ML approaches often rely on idealized simulations and limited attack scenarios. They tend to omit realistic conditions such as channel loss, phase noise, photon shot noise, depolarization, and detector imperfections, and often ignore critical quantum-level threats such as PNS, Intercept-and-Resend, Trojan Horse, RNG, Detector Blinding, and Wavelength-Dependent attacks. Furthermore, to the best of our knowledge, Quantum Machine Learning has yet to be applied to QKD intrusion detection.
To bridge this gap, we propose a hybrid intrusion detection system that integrates Quantum Long Short-Term Memory (QLSTM) into QKD monitoring. By combining quantum-enhanced learning with classical deep learning, the proposed model aims to improve detection accuracy and adaptability to evolving threats. We also introduce a new simulated dataset for decoy-state BB84 DV-QKD under eight scenarios: normal operation, Intercept-and-Resend, Trojan Horse, PNS, RNG, Wavelength-dependent Trojan Horse, Detector Blinding, and combined attacks.
I-A Contributions
The contributions of this work can be summarized as follows:
  • Comperhenisve QKD dataset: Given the lack of publicly available and comprehensive datasets, this work proposes a realistic QKD simulation using the PennyLane library. The dataset simulates decoy-state BB84 DVQKD under eight different scenarios: normal QKD operation, Intercept-and-Resend attack, Trojan Horse attack, PNS attack, RNG attack, Wavelength-dependent Trojan Horse attack, Detector Blinding attack, and combined attack. It also includes various quantum metrics such as Quantum Bit Error Rate (QBER), measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring a dataset that accurately reflects real-world conditions.
  • Quantum Machine Learning for quantum intrusion detection on QKD systems: A hybrid intrusion detection model is proposed, integrating a QLSTM network with a classical LSTM. This approach leverages the advantages of both quantum computation and classical deep learning, offering a practical intrusion detection solution for future QKD networks.
I-B Organization
The rest of this paper is organized as follows. Section II introduces the proposed Hybrid QLSTM model for detecting QKD attacks. Section III discusses the dataset generation process including the simulation algorithm that was used to generate the dataset. Section IV illustrates the evaluation setup and the used metrics. Section V presents and analyzes the experimental results, focusing on the performance of the proposed hybrid QLSTM model. Finally, Section LABEL:sec:Section_7 concludes this paper and provides suggestions for future work.
II Model
This section presents the Hybrid QLSTM Model for QKD attack detection. Given the sequential nature of QKD transmissions, QLSTM is used to analyze quantum security metrics over time, improving attack detection accuracy compared to classical LSTM. The model consists of a QLSTM layer, a classical LSTM layer, and a fully connected layer. The QLSTM captures quantum-enhanced sequential features, the LSTM refines temporal dependencies, and the final layer maps features to logits for classification. The chapter also covers GPU-accelerated training, using CrossEntropy Loss, AdamW optimization, and a cosine annealing learning rate schedule. The model is trained for 50 epochs, with early stopping to prevent overfitting.
In this work, we use QLSTM to analyze QKD security metrics over time and detect anomalies. QLSTM is an extension of the classical Long Short-Term Memory (LSTM) model, incorporating quantum principles, to enhance learning efficiency in sequential data tasks. Traditional LSTMs are widely used in time-series analysis, making them well-suited for detecting patterns in QKD-based attack scenarios. However, QLSTM leverages quantum mechanics to capture dependencies more effectively than classical models. As shown in [17], QLSTM outperformed classical LSTMs in detecting anomalies in Distributed Denial of Service (DDoS) attacks. Given the sequential nature of QKD transmissions, detecting anomalies in quantum metrics is crucial. Compared to classical LSTM, QLSTM converges faster and achieves higher accuracy in detecting attack patterns, making it a strong candidate for quantum security applications.
The model consists of three main layers: a QLSTM layer, a classical LSTM layer, and a fully connected layer. Figure 1 depicts the high-level design of the proposed model. The QLSTM layer processes the sequential input data while incorporating quantum properties, such as entanglement and superposition, capturing complex features in the data. The output of the QLSTM layer is fed into a classical LSTM layer, which learns the temporal patterns from the quantum-enhanced sequences. Finally, a fully connected layer maps the learned features to the output space, producing logits, which are the scores that serve as the output of an ML model. These logits are passed to a softmax function for classification. Pseudocode 1 describes the model.
Figure 1: High Level Diagram of The Proposed Hybrid QLSTM Model
Algorithm 1 Hybrid QLSTM Model
1:1. Quantum LSTM (QLSTM) Layer Initialization
2:Set parameters:
  • input_sizeinput\_size
  • hidden_sizehidden\_size
  • n_qubits=9n\_qubits=9
  • n_qlayers=1n\_qlayers=1
  • $backend=$ “default.qubit”
3:Use Strongly Entangling Layers for quantum gates
4:Initialize quantum circuit weights and entanglement parameters
5:Define quantum gates for input, forget, output, and candidate cell states
6:Post-process quantum measurements for classical compatibility
7:2. Classical LSTM Layer Initialization
8:Set input size =n_qubits=9=n\_qubits=9
9:Set hidden size =32=32
10:Initialize weights and hidden/cell states
11:3. Fully Connected (FC) Layer
12:Define linear layer mapping $32\rightarrow output\_dim=4$
13:4. Forward Pass
14:function Forward($x$)
15:  Input: $x$ of shape (batch_size,seq_len,feature_dim)(batch\_size,seq\_len,feature\_dim)
16:  Initialize $h_{0},c_{0}$ for QLSTM as zeros
17:  for each time step $t$ in seq_lenseq\_len do
18:   $x_{t}\leftarrow x[:,t,:]$ $\triangleright$ Extract features at time $t$
19:   $h_{t},c_{t}\leftarrow\text{QLSTM}(x_{t},h_{t-1},c_{t-1})$ $\triangleright$ Quantum processing
20:   Store $h_{t}$
21:  end for
22:  $qlstm\_output\leftarrow$ Stack all $h_{t}$ along time axis
23:  $lstm\_output,(h_{n},c_{n})\leftarrow\text{ClassicalLSTM}(qlstm\_output)$
24:  $last\_hidden\leftarrow h_{n}$ $\triangleright$ Or lstm_output[:,1,:]lstm\_output[:,-1,:]
25:  $logits\leftarrow\text{FC}(last\_hidden)$
26:  return logits
27:end function
Figure 2: (a) QLSTM and (b) LSTM Model Architectures
II-A QLSTM Layer
This layer is implemented based on the open-source QLSTM implementation in [18]. However, for the purpose of this project, the original entanglement layer was changed from Basic Entanglement to Strongly Entanglement Layer. At each time step, the QLSTM receives the current input $x_{t}$ and the previous hidden and cell states, $h_{t}$ and $c_{t}$, respectively. It processes this information through variational quantum circuits (VQCs), representing the Forget Gate, Input Gate, Candidate Cell State, and Output Gate. Specifically, the gates are computed as follows:
$i_{t}$ $=\sigma(VQC_{2}(v_{t}))\quad f_{t}=\sigma(VQC_{1}(v_{t})),$ (1)
$\tilde{C}_{t}$ $=\tanh(VQC_{3}(v_{t})),\quad o_{t}=\sigma(VQC_{4}(v_{t})),$ (2)
where, $v_{t}$ is the $N$-dimensional input vector at time step $t$, $VQC_{k}(\cdot)$ represents a parameterized variational quantum circuit for gate $k$, $\sigma$ denotes the sigmoid activation function that regulates information flow, $\tanh$ denotes the hyperbolic tangent activation function for candidate states, and $i_{t},f_{t},\tilde{C}_{t},o_{t}$ are the input, forget, candidate cell state, and output gates at time $t$, respectively.
This information is processed through variational quantum circuits (VQCs), Forget Gate, Input Gate, Update Gate, and Output Gate. The outputs of the VQCs are passed through classical nonlinearities ($\sigma$, $\tanh$) before combining to produce the updated hidden and cell states. These updated states are then propagated to the next time step. The architecture of the QLSTM model used in this work is visually illustrated in Figure 2 (a).
II-B Classical LSTM Layer
This layer consists of 32 hidden states and operates in three main steps:
Input Processing
The quantum layer’s output is reshaped into a sequence format to match the LSTM’s expected input $x_{t}$.
LSTM Processing
At each time step $t$, the LSTM layer updates its internal memory using the input $x_{t}$ and the previous hidden and cell states ($h_{t-1}$, $c_{t-1}$). As illustrated in Figure 2 (b), this process involves three gates, the forget gate $f_{t}$, input gate $i_{t}$, and output gate $o_{t}$, along with a candidate cell state $\tilde{c}_{t}$. The gate activations are computed as follows:
$i_{t}$ $=\sigma\big{(}W_{xi}x_{t}+W_{hi}h_{t-1}+b_{i}\big{)},$ (3)
$f_{t}$ $=\sigma\big{(}W_{xf}x_{t}+W_{hf}h_{t-1}+b_{f}\big{)},$ (4)
$\tilde{c}_{t}$ $=\tanh\big{(}W_{xc}x_{t}+W_{hc}h_{t-1}+b_{c}\big{)},$ (5)
$o_{t}$ $=\sigma\big{(}W_{xo}x_{t}+W_{ho}h_{t-1}+b_{o}\big{)}.$ (6)
where, $\sigma(\cdot)$ is the sigmoid activation function, $\tanh(\cdot)$ is the hyperbolic tangent activation function, $x_{t}$ is the current input vector, $h_{t-1}$ and $c_{t-1}$ are the hidden and cell states from the previous time step, $i_{t}$, $f_{t}$, and $o_{t}$ are the input, forget, and output gate activations, respectively, $\tilde{c}_{t}$ is the candidate cell state, $W_{x\ast}$ and $W_{h\ast}$ are the learnable weight matrices, and $b_{\ast}$ are the bias vectors for each gate.
Hidden State Extraction
Once the sequence is fully processed, the last hidden state $h_{t}$ of the LSTM model is extracted to be used in the final prediction step.
$c_{t}=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}_{t},$ (7)
where $\odot$ denotes element-wise multiplication. Finally, the hidden state $h_{t}$ is updated as:
$h_{t}=o_{t}\odot\tanh(c_{t}).$ (8)
This ensures that the output at each step is a filtered representation of the current cell state. After processing the entire sequence, the final hidden state $h_{t}$ is extracted and passed into the next stage for classification.
II-C Fully Connected Layer
The fully connected layer maps output from the classical LSTM into the final output space, which consists of 4 logits, each corresponding to a possible class (Normal, Intercept-and-Resend Attack, PNS Attack, Trojan-Horse Attack). These logits are later transformed to class probabilities using a softmax function during evaluation.
III Dataset
This Section introduces a realistic QKD dataset designed to address the limitations of existing datasets, which often rely on mathematical assumptions or lack specific attack scenarios. By incorporating PNS, intercept-and-resend, Trojan-Horse, RNG, detector blinding, wavelength dependent, and combined attacks, the dataset better represents real-world threats to quantum communication. Key quantum metrics are collected to evaluate the security of the QKD process. The dataset goes through preprocessing steps, including feature selection, normalization, and augmentation, to ensure its suitability for the proposed hybrid QLSTM model. This dataset serves as a benchmark for evaluating the effectiveness of the QLSTM model in detecting QKD security threats. Existing QKD datasets (as discussed in Section I) often suffer from significant limitations, such as containing very few features, relying on mathematical assumptions, or focusing on the practical vulnerabilities of QKD systems. In addition, many studies focus on general eavesdropping threats without specifying the exact attack strategy or emphasizing the imperfections of the device. In contrast, this work explicitly incorporates PNS, Intercept-and-Resend, and Trojan-Horse attacks, which directly exploit QKD vulnerabilities at the photon transmission level. The selection of these attacks is driven by gaps identified in previous research. By addressing these threats, our dataset enables a more precise evaluation of ML-based detection mechanisms.
In this work, we adopt decoy-state BB84 protocol, which enhances security by randomly transmitting additional photon (decoy) pulses with varying intensities. Alice randomly sends both signal and decoy pulses, and Bob measures them as in standard BB84. Since Eve cannot distinguish between signal and decoy pulses, she cannot selectively measure only multi-photon pulses without being detected. By analyzing detection rates for both signal and decoy states, Alice and Bob can identify and mitigate PNS attacks before finalizing the key [19].
The simulation of the decoy-state BB84 protocol includes four scenarios: normal QKD operation, Intercept-and-Resend attacks, Trojan-Horse attacks, and Photon-Number-Splitting (PNS) attacks. It collects various quantum metrics, including the Quantum Bit Error Rate (QBER), bit mismatch entropy, signal and decoy detection rates, and time-based statistics, discussed in table V. The simulation is carried out to generate approximately a 256-bit secret key while reflecting practical transmission conditions. To make the dataset more realistic, different noise sources are incorporated. In particular, quantum channel losses are introduced to simulate photon attenuation over the link, photon shot noise is added to reflect the inherent fluctuations in photon detection, and depolarizing noise is applied to represent random state decoherence that naturally occurs during transmission. These noise sources vary randomly for each photon transmission, with their strength set to low, moderate, or high levels. While this does not capture every possible imperfection in a real system, it provides a closer approximation to realistic conditions compared to purely ideal simulations. Table III presents the distribution of attack scenarios in the dataset, which consists of 10,329 samples evenly spread across eight labels. In the following subsections, we describe each scenario and give details on how the effect of each condition was taken into account as we develop the dataset. Table IV summarizes the conditions we considered in each scenario.
TABLE III: Distribution of QKD Attack Scenarios
Scenario Samples
Normal QKD 1292
Man-in-the-Middle (MITM) 1291
Photon Number Splitting (PNS) 1291
Trojan-Horse Attack 1291
Wavelength Dependent Trojan-Horse 1291
Random Number Generator Attack 1291
Detector Blinding Attack 1291
Combined Attack 1291
TABLE IV: Summary of Attack Scenarios and Conditions Considered
Scenario Description Key Conditions Considered Expected Impact on QKD
Normal QKD Standard QKD process without any attack. Alice and Bob exchange a key securely. No external interference, stable noise levels, and proper key sifting. Low QBER, stable detection rates, and normal transmission times.
Intercept-Resend Attack Eve intercepts and measures photons before resending them to Bob. Eve randomly measures in a basis, causing errors in transmission. Increased QBER, mismatch entropy, and detection errors.
Trojan-Horse Attack Eve injects additional photons into Alice’s system to extract internal settings. Eve controls the intensity and timing of injected photons. Potential bias in key generation and increased information leakage.
Photon-Number Splitting (PNS) Attack Eve selectively measures multi-photon pulses and retains a copy of the key. Eve detects photon pulses without altering single-photon transmissions. Lower detection rates without a significant increase in QBER.
Wavelength-Dependent Trojan-Horse Attack Eve injects pulses at off-wavelengths to exploit detector sensitivity. Wavelength differences affect detection efficiency and phase encoding. Increased QBER, timing delays, and reduced detection efficiency.
Random Number Generator (RNG) Attack Eve exploits bias or predictability in Alice’s and Bob’s random number generators. Biased or predictable bit and basis choices due to compromised RNGs. Increased predictability of key bits, subtle bias in entropy, possible QBER changes.
Detector Blinding Attack Eve blinds Bob’s detectors with bright light and controls their outputs. Detectors forced into classical mode, responding only to Eve’s signals. Lower detection efficiency, possible QBER increase, key fully compromised if undetected.
Combined Attack Eve combines multiple strategies (e.g., RNG, detector blinding, wavelength injection). Simultaneous exploitation of multiple vulnerabilities. Mixed effects: increased QBER and entropy, reduced efficiency, and masked attack traces.
TABLE V: Summary of Quantum Key Distribution (QKD) Metrics and Corresponding Functions
Metric Description Formula
Key Length The length of the final generated secret key. N/A
QBER The ratio of mismatched bits after the sifting process to the total sifted key bits. Indicates the level of errors in the QKD process. $QBER=\frac{\text{total mismatches}}{\text{total sifted bits}}$
Measurement Entropy Shannon entropy for matches and mismatches, measuring uncertainty in the outcomes [20]. $H=-\sum p_{i}\log_{2}p_{i}$
Signal Detection Rate Ratio of detected signal pulses to total sent signal pulses. $\frac{\text{signal detections}}{\text{sent signal pulses}}$
Decoy Detection Rate Ratio of detected decoy pulses to total sent decoy pulses. $\frac{\text{decoy detections}}{\text{sent decoy pulses}}$
Signal Loss Rate Ratio of lost signal pulses to total sent signal pulses. $\frac{\text{lost signal pulses}}{\text{sent signal pulses}}$
Decoy Loss Rate Ratio of lost decoy pulses to total sent decoy pulses. $\frac{\text{lost decoy pulses}}{\text{sent decoy pulses}}$
Average Photon Transmission Time Average time taken for a photon to be transmitted and detected. $\text{Average Time}=\frac{1}{N}\sum_{i=1}^{N}T_{i}$
Whole Key Transmission Time Total time taken for all photons in the key exchange process. $\sum T_{i}$
Arrival Time Variance Variance in the time taken for photon transmission, measuring fluctuations. $\sigma^{2}=\frac{1}{N}\sum_{i=1}^{N}(T_{i}-\bar{T})^{2}$
Arrival Time Deviation Mean absolute deviation from the average photon transmission time. $\frac{1}{N}\sum_{i=1}^{N}|T_{i}-\bar{T}|$
III-A Attack Scenarios
Normal QKD Operation
In a normal DV-QKD scenario, where there is no presence of Eve, the process of exchanging the key follows the standard BB84 protocol, where Alice randomly prepares qubits in rectilinear (horizontal/vertical) or diagonal ($\pm 45^{\circ}$) bases and Bob randomly chooses a basis for measurement. After transmission, they publicly compare their chosen bases and keep only the bits where their bases match (sifting). Under these conditions, the QBER and Measurement Entropy are as close to zero as possible, indicating minimal errors and maximum predictability in the measurement outcomes. Moreover, the Signal and Decoy Detection Rates should be as close to the predefined distribution ratio of signal and decoy pulses as possible, showing that no external interference occurs. Additionally, the time-based metrics should remain stable and consistent. Any deviations from these expected values could suggest the presence of noise, losses, or an external attack on the quantum channel. Algorithm 2 illustrates the pseudo-code for the normal QKD simulation.
Algorithm 2 Normal QKD Simulation
1:Initialize: Sifted key $K=[]$; counters $N_{\text{sig}},N_{\text{dec}},N_{\text{loss,sig}},N_{\text{loss,dec}},N_{\text{err}},N_{\text{det,sig}},N_{\text{det,dec}}=0$; time lists $T_{\text{tx}},T_{\text{rx}}=[]$
2:for $i=1$ to $N_{\text{trans}}$ do $\triangleright$ $N_{\text{trans}}=50$
3:  Randomly assign pulse type: signal ($p_{\text{sig}}=0.7$) or decoy ($p_{\text{dec}}=0.3$)
4:  Simulate noise and photon loss
5:  if photon is lost then
6:   Increment $N_{\text{loss,sig}}$ or $N_{\text{loss,dec}}$
7:   continue
8:  end if
9:  Alice selects random bit $a\in\{0,1\}$ and basis $b_{A}\in\{0,1\}$
10:  Encode qubit as $|\psi\rangle=\text{Encode}(a,b_{A})$
11:  Bob randomly selects measurement basis $b_{B}\in\{0,1\}$
12:  Bob measures: $b=\text{Measure}(|\psi\rangle,b_{B})$
13:  Apply bit flip error: $b\leftarrow b\oplus\text{Bern}(p_{\text{err}})$
14:  Apply depolarization and photon shot noise to $b$
15:  $b\leftarrow\begin{cases}b,&\text{with probability }1-p_{\text{depol}}\\ \text{random}(0,1),&\text{with probability }p_{\text{depol}}\end{cases}$
16:  Apply photon shot noise using the Poisson model: $P(k|\bar{n})=\dfrac{\bar{n}^{k}e^{-\bar{n}}}{k!}$, where $k$ is the number of detected photons and $\bar{n}$ is the mean photon number
17:  Record $T_{\text{tx}}[i],T_{\text{rx}}[i]$
18:  if $b_{A}=b_{B}$ then
19:   Append $(a,b)$ to $K$ (sifted key)
20:   if $a=b$ then
21:     Increment $N_{\text{det,sig}}$ or $N_{\text{det,dec}}$
22:   else
23:     Increment $N_{\text{err}}$
24:   end if
25:  end if
26:end for
[…]
32:Return: Sifted key $K$, QBER, entropy $H$, $R_{\text{sig}}$, $R_{\text{dec}}$, $L_{\text{sig}}$, $L_{\text{dec}}$, time metrics $\Delta T$
Intercept-and-Resend Attack
In the Intercept-and-Resend attack scenario, the QKD process is intercepted by Eve, who measures the transmitted photons and then retransmits them to Bob, effectively impersonating Alice. This process typically introduces detectable errors, but if Eve only attacks a small fraction of signals, the errors might remain below the detection threshold. Under these conditions, the QBER and Measurement Entropy significantly increase, indicating potential errors and greater unpredictability in the measurement outcomes. Furthermore, the signal and decoy detection rates should be reduced due to Eve already measuring the signal or the decoy, which leads to possible external interference or tampering. Additionally, the time-based metrics are expected to increase because of the delay introduced by Eve’s interception. Based on these assumptions, the simulated attack is constructed as described in Algorithm 3.
Algorithm 3 Intercept-and-Resend Attack Simulation
1:Follows Normal QKD States 1-10.
2:Eve intercepts the photon:
3:  Eve randomly selects a basis $b_{E}\in\{0,1\}$
4:  Eve measures the photon: $e=\text{Measure}(|\psi\rangle,b_{E})$
5:  Apply bit flip error to Eve’s measurement: $e\leftarrow e\oplus\text{Bern}(p_{\text{err,Eve}})$
6:Apply depolarization noise:
$e\leftarrow\begin{cases}e,&\text{with probability }1-p_{\text{depol,Eve}}\\ \text{random}(0,1),&\text{with probability }p_{\text{depol,Eve}}\end{cases}$
7:  Eve prepares a new qubit $|\psi^{\prime}\rangle=\text{Encode}(e,b_{E})$ and sends it to Bob
8:Follows Normal QKD States 9-32
PNS Attack
In the PNS attack scenario, Eve intercepts and stores a single photon from multi-photon pulses while allowing the remaining photons to reach Bob. In this attack, practical QKD systems that rely on weak coherent pulses can emit multi-photon states, which makes them vulnerable to selective photon extraction by Eve, which primarily affects the detection rates, causing a noticeable decrease. However, since Eve does not perform any measurements on the intercepted photons, the QBER and Measurement Entropy remain unchanged, as no additional errors or randomness are introduced into the quantum channel. Furthermore, the time-based metrics remain relatively stable, as the interception process does not introduce significant delays in the transmission.
Algorithm 4 PNS Attack Simulation
1:Follows Normal QKD States 1-9.
2:Simulate photon intensity:
  • Generate a random number $r\in[0,1]$
  • If $r<0.8$, set pulse type to single-photon
  • Else, set pulse type to multi-photon
3:Eve’s Photon Number Splitting (PNS) Attack:
4:Eve checks if the pulse is multi-photon.
5:if multi-photon pulse then
6:  Eve intercepts and stores one photon.
7:  Allows remaining photons to pass without measurement.
8:else
9:  Eve does nothing.
10:end if
11:Follows Normal QKD States 9-32
Trojan Horse Attack
In the Trojan horse attack scenario, Eve injects additional photons into Alice’s system by sending bright light pulses into the QKD device on Alice’s or Bob’s side to extract internal information such as basis choice or key bits [21]. This attack primarily affects the detection rates, leading to an increase as more photons are detected on Bob’s side. However, since Eve does not perform direct measurements on the transmitted photons, the QBER and Measurement Entropy do not increase significantly, with mismatches only occurring if Bob measures an injected pulse. In addition, the time-based metrics remain stable, as the attack does not introduce notable transmission delays. Researchers in [22] indicated that Eve can adjust the photon injection strategy by varying between strong and weak pulses based on the detected reflection patterns. Based on these assumptions, the simulated attack is constructed as described in Algorithm 5.
Algorithm 5 Trojan-Horse Attack Simulation
1:Follows Normal QKD States 1-10.
2:Eve’s Trojan-Horse Attack:
3:Generate a random number $r_{1}\in[0,1]$
4:If $r_{1}<p_{\text{inject}}$ (injection occurs):
5:  Generate another random number $r_{2}\in[0,1]$
6:If $r_{2}<0.5$ (strong pulse):
7:   Set Bob’s detection probability to high (e.g., $p_{\text{detect,strong}}$)
8:   Generate $r_{3}\in[0,1]$; if $r_{3}<0.5$, set detected signal as mismatched
9:Else (weak pulse):
10:   Set Bob’s detection probability to low (e.g., $p_{\text{detect,weak}}$)
11:   Generate $r_{3}\in[0,1]$; if $r_{3}<0.2$, set detected signal as mismatched
12:  Continue as normal QKD process for this pulse
13:Else: No injection; proceed as normal QKD process
14:Follows Normal QKD States 9-32
Wavelength Dependent Trojan-Horse Attack
In Wavelength Dependent Trojan-Horse Attack, Eve exploits the wavelength sensitivity of QKD components by injecting bright pulses at various wavelengths, rather than only at the legitimate signal wavelength [23, 22]. This approach leverages the fact that the detection efficiency and phase encoding in optical devices can vary significantly with wavelength. By carefully selecting the injected wavelength, Eve can reduce detection efficiency, introduce phase errors, and increase her chances of eavesdropping without being detected. Unlike the classic Trojan-Horse Attack, this variant can also affect QBER and timing metrics due to wavelength-dependent delays and errors. Researchers have shown that such attacks may introduce subtle but measurable changes in the QKD statistics, potentially enabling Eve to gain information about the key while minimizing her detection risk [22, 21].
Based on these considerations, the simulated attack is constructed as described in Algorithm 6.
Algorithm 6 Wavelength Dependent Trojan-Horse Attack Simulation
1:Follows Normal QKD States 1-10.
2:Eve’s Wavelength-Dependent Trojan-Horse Attack:
3:Generate a random number $r_{1}\in[0,1]$
4:If $r_{1}<p_{\text{inject}}$ (injection occurs):
5:  Randomly select attack wavelength $\lambda_{\text{attack}}$ from possible set
6:  Calculate wavelength difference $\Delta\lambda=|\lambda_{\text{attack}}-\lambda_{\text{legit}}|$
7:  Update detection efficiency: $\eta\leftarrow\eta_{0}\cdot f(\Delta\lambda)$
8:  Update phase encoding: $\theta\leftarrow\theta+\phi(\Delta\lambda)$
9:  Generate $r_{2}\in[0,1]$
10:If $r_{2}<p_{\text{eavesdrop}}$:
11:   Eve successfully extracts information about the key bit
12:If $\Delta\lambda>0$:
13:   With probability $p_{\text{error}}$, introduce additional measurement error at Bob’s side
14:  Add wavelength-dependent delay to photon arrival time
15:  Continue as normal QKD process for this pulse
16:Else: No injection; proceed as normal QKD process
17:Follows Normal QKD States 9-32
Random Number Generator Attack
The Random Number Generator (RNG) Attack targets the fundamental assumption that Alice and Bob’s random choices of bits and bases are truly unpredictable and uniformly distributed. If the RNGs used in the QKD protocol are biased, patterned, or otherwise predictable—due to hardware flaws or deliberate compromise—an eavesdropper (Eve) can exploit this to increase her information about the key [24, 25]. In this attack, Eve analyzes recent output patterns from the RNG and leverages statistical bias or repeated patterns to predict future choices with non-negligible probability. If Eve can correctly guess Alice’s bit and basis, she can intercept and resend the photon, introducing minimal disturbance and remaining largely undetected. The effectiveness of the attack depends on the degree of bias, the pattern predictability, and Eve’s ability to exploit them in real time. Algorithm 7 outlines the simulation process for this attack.
Algorithm 7 RNG Attack Simulation
1:Follows Normal QKD States 1-10.
2:Set bit bias $b_{\text{bit}}$, basis bias $b_{\text{basis}}$, pattern exploitation probability $p_{\text{pattern}}$, prediction window size $w$
3:for each photon transmission do
4:  Alice generates bit and basis using compromised RNG with bias and pattern memory
5:  Record previous bits and bases for pattern prediction
6:  If Eve predicts bit and basis (using recent history and $p_{\text{pattern}}$):
7:    With probability $p_{\text{intercept}}$, Eve intercepts and resends photon
8:    If Eve’s prediction is correct and bases match, she may introduce errors
9:  Simulate transmission noise and photon loss
10:  Bob chooses basis using similarly compromised RNG
11:  Bob measures received photon and records result
12:end for
13:Sift key: keep only cases where both Alice and Bob have valid detections and matching bases
14:Calculate QBER, detection/loss rates, bias metrics, and prediction success rate
Detector Blinding Attack
In Detector Blinding Attack, Eve manipulates the single-photon detectors at Bob’s side by sending bright continuous-wave light, forcing them into a linear (classical) mode where they no longer behave as true quantum detectors [26]. While blinded, Bob’s detectors lose their single-photon sensitivity and only respond to Eve’s specifically tailored trigger pulses, allowing Eve to control which detector clicks and thus gain full knowledge of the key without introducing detectable errors. The attack is typically performed in bursts (blinding duration), and Eve can inject her own signals with high efficiency while the detectors are blinded. During these periods, the detection efficiency drops, and the quantum bit error rate (QBER) may increase if the attack is not perfectly synchronized. However, if performed carefully, the attack can leave the QBER and other statistics nearly unchanged, making detection extremely challenging [26, 27]. Algorithm 8 summarizes the simulation process for this attack.
Algorithm 8 Detector Blinding Attack Simulation
1:Follows Normal QKD States 1-10.
2:Set blinding probability $p_{\text{blind}}$, blinding duration $d_{\text{blind}}$, injection efficiency $p_{\text{inject}}$
3:Initialize detector state as normal with efficiency $\eta_{\text{normal}}$
4:for each photon transmission do
5:  Randomly select photon as signal (70%) or decoy (30%)
6:  If Eve blinds detectors ($r_{1}<p_{\text{blind}}$ and not already blinded):
7:    Set detector state to blinded for $d_{\text{blind}}$ rounds
8:    Set efficiency to $\eta_{\text{blinded}}\ll\eta_{\text{normal}}$
9:  Alice prepares qubit (bit, basis), applies encoding
10:  Simulate transmission noise and losses
11:  If photon lost: record as loss, continue
12:  If detectors are blinded:
13:    With probability $p_{\text{inject}}$, Eve injects her own signal and chooses basis/bit
14:    Use Eve’s parameters for measurement
15:  Bob measures qubit in random basis
16:  Simulate detection based on current detector efficiency
17:  If detection occurs:
18:    Bob records measured bit
19:  Else: record as loss
20:  Update blinding countdown; if expired, restore detector to normal
21:end for
22:Sift key: keep only matching bases and successful detections
23:Calculate QBER, detection/loss rates, and timing metrics
Combined Attack
The Combined Attack represents a realistic and formidable threat scenario in which an eavesdropper (Eve) employs multiple quantum hacking strategies simultaneously to maximize her chances of compromising the QKD system. In this attack, Eve leverages a combination of wavelength attacks, detector blinding, and random number generator (RNG) manipulation. By exploiting device imperfections and protocol vulnerabilities in concert, Eve can bypass optical filters with off-wavelength photons [28], force detectors into a controllable linear mode [26], and predict or bias random choices made by Alice and Bob [25]. The synergy between these attacks allows Eve to increase her information gain while minimizing detection, as the errors and losses introduced by one attack may mask or compensate for those caused by another. Algorithm 9 summarizes the simulation process for this multifaceted attack.
Algorithm 9 Combined Attack Simulation
1:Follows Normal QKD States 1-10.
2:Set probabilities for wavelength attack, detector blinding, and RNG attack
3:for each photon transmission do
4:  Randomly activate any subset of the three attacks per photon
5:  Simulate photon loss and noise
6:  Alice prepares bit and basis (biased/predictable if RNG attack active)
7:  Eve may intercept if she predicts Alice’s choices (RNG attack)
8:  If wavelength attack active, modify photon efficiency and phase
9:  If detector blinding active, manipulate Bob’s detection probability and error rate
10:  Bob chooses basis (biased if RNG attack active) and measures photon
11:  Apply measurement errors from all active attacks
12:end for
13:Sift key: keep only cases with valid detections and matching bases
14:Calculate QBER, detection/loss rates, attack success rates, and timing metrics
III-B Generation
Our dataset is constructed by executing multiple iterations of the QKD simulations discussed in the previous subsections. Each iteration simulates all eight possible scenarios (normal QKD process, Intercept-and-Resend attack, PNS attack, Trojan horse attack, wavelength-dependent Trojan horse, RNG attack, detection blinding attack, and combined attack), capturing a set of QKD metrics. These include quantum bit error rate (QBER), measurement entropy, signal and decay detection rates, signal and decay loss rates, photon transmission time metrics, and time-based variation metrics. These metrics provide insights into the security of the quantum channel and help detect anomalies in the process. A total of 1292 iterations per scenario were performed, resulting in a balanced dataset of 10,329 samples. Table VI shows how the first few lines of the proposed dataset. Data preprocessing is a crucial step in the development of any ML or DL model, as it significantly affects the model’s accuracy and performance. In this context, the preprocessing framework for the proposed model contains several steps to prepare the dataset for training and evaluation. These steps include feature selection, feature transformation, dataset splitting, noise injection, and input representation.
TABLE VI: Sample of the Quantum Key Distribution dataset across 8 attack scenarios
Key_Length QBER Measurement_entropy Signal_detection_rate Decoy_detection_rate Signal_loss_rate Decoy_loss_rate Avg_Photon_time Whole_key_time Arrival_var Arrival_dev Label
253 0.0553 0.2981 0.7311 0.7545 0.2691 0.2455 0.0684 35.34 0.0019 0.0359 normal
247 0.3927 0.9661 0.7510 0.7110 0.2489 0.2889 0.1879 97.13 0.0213 0.1059 mitm_attack
284 0.0739 0.3798 0.7112 0.6446 0.2442 0.3013 0.0923 47.62 0.0059 0.0653 pns_attack
240 0.2542 0.8169 0.7437 0.7679 0.2563 0.2321 0.0932 49.02 0.0071 0.0672 trojan_horse_attack
311 0.0129 0.0990 0.8663 0.8738 0.1337 0.1262 0.0630 44.12 0.0024 0.0380 wavelength_dependent_trojan_attack
454 0.1850 0.6882 0.9090 0.8943 0.0901 0.1057 0.0697 48.79 0.0040 0.0479 rng_attack
142 0.0845 0.4178 0.4340 0.4265 0.2740 2.6567 0.0581 40.69 0.0096 0.0734 detector_blinding_attack
344 0.1744 0.6585 0.8198 0.8815 0.1820 0.1185 0.0653 45.69 0.0025 0.0374 combined_attack
Feature Selection
The dataset contains 11 features, but only 9 features have been selected for training the proposed model, excluding the Signal Detection Rate and Decoy Detection Rate. This decision is based on the fact that the Signal Loss Rate and Decoy Loss Rate are already included, and they imply the same information. Removing these features helps prevent overfitting and ensures that the model generalizes well.
Feature Transformation
All feature values are standardized using Z-score normalization (StandardScaler) to ensure that they have a mean of zero and a standard deviation of one, using the following equation:
$X^{\prime}=\frac{X-\mu}{\sigma}$ (9)
where $\mu$ represents the mean of the feature values, and $\sigma$ denotes the standard deviation. Additionally, the categorical labels are encoded using the one-hot encoding technique to ensure that the model correctly interprets the categorical class information.
Dataset Splitting
The dataset is split into training and testing sets with an 80%-20% ratio to ensure sufficient training data while maintaining a reliable evaluation set.
Noise Injection
Gaussian noise is added to the training set to improve model robustness and enhance the generalization capability. This noise follows a normal distribution:
$X_{\text{train noisy}}=X_{\text{train}}+\mathcal{N}(\mu,\sigma^{2})$ (10)
where $\mathcal{N}(\mu,\sigma^{2})$ represents Gaussian noise with mean $\mu=0$ and standard deviation $\sigma=0.05$. This ensures that the model is exposed to slightly varied training inputs, making it more resilient to real-world variations.
III-B1 Input Representation
The input batch size is set to 64 samples per batch, ensuring a balance between training efficiency and model stability. The sequence length is fixed at 9 timesteps, corresponding to the 9 selected features extracted from the QKD system.
IV Evaluation
This section discusses the evaluation criteria and presents the performance results of the proposed hybrid QLSTM model compared to classical CNN and LSTM models, followed by a comparative analysis with existing work in the literature. All models were evaluated using the dataset discussed in Section III-B, after applying the same preprocessing steps detailed in the same section, and trained using the same setup described in Section IV-A to ensure fair evaluation.
IV-A Setup
The three architectures QLSTM, LSTM, and CNN were trained under identical conditions. Training was conducted on a GPU (Google Colab); both model parameters and data tensors were moved to the GPU device to leverage accelerated matrix operations and, in the case of the QLSTM, faster quantum‐circuit simulations. We used PyTorch’s CrossEntropyLoss, which applies a softmax over the logits followed by the multi‐class log‐loss. Optimization was performed using the AdamW optimizer with initial learning rate $5\times 10^{-4}$ and weight decay $1\times 10^{-4}$. A Cosine Annealing Warm Restarts scheduler with restart period $T_{0}=50$ epochs was applied to cyclically modulate the learning rate. Mini‐batches of 64 samples were used throughout the training. Early stopping with patience of 5 epochs monitored the validation loss, halting training if no improvement was observed for five consecutive epochs and thus avoiding overfitting. Each model was trained for 10, 20, and 50 epochs under this same protocol. These parameters are summarized in Table VII.
TABLE VII: Hyperparameters Used for QLSTM Training
Hyperparameter Value
Optimizer AdamW
Learning rate 0.005
Weight decay $1\times 10^{-4}$
Scheduler
Cosine Annealing
Warm Restarts ($T_{0}=50$)
Batch size 64
Early stopping patience 5
IV-B Metrics
The model is evaluated based on multiple performance metrics, including accuracy, F1-score, recall, and precision, to assess its ability to detect attacks on a QKD system. These metrics are computed using the following formulas:
Accuracy
Represents the overall correctness of the model by measuring the proportion of all predictions (both attack and normal) that are correctly classified.
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (11)
where TP (True Positives) is the number of attack labels that are correctly classified as attacks, TN (True Negatives) is the number of normal labels that are correctly classified as normal, FP (False Positives) is the number of normal labels that are classified as attacks, and FN (False Negatives) is the number of attack labels that are classified as normal.
Precision
Measures how many predicted attacks are actual attacks, reducing false positives.
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (12)
Recall
Evaluates how well the model detects actual attacks, minimizing false negatives.
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (13)
F1-score
Balances both precision and recall, useful when there is an imbalance between normal and attack labels.
$\text{F1-score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (14)
Furthermore, the model’s performance is compared against classical CNN and LSTM models, all trained under the same conditions, to assess the proposed model’s effectiveness on multi-class classification. Although CNN model is traditionally designed for spatial data, in this study, we adapted it using 1D convolutional layers by reshaping the input such that the 9 features form a one-dimensional structure.
V Results
The evaluation results for all models demonstrate that the proposed QLSTM model outperforms both LSTM and CNN in detecting quantum attacks on the QKD system. Table VIII shows that even in the early stages of training (10 epochs), QLSTM achieves an accuracy of 88.3%, already higher than LSTM (85.8%) and CNN (86.8%). This early lead reflects QLSTM’s ability to capture temporal and statistical patterns present in quantum attack traffic, which are more challenging for classical models to extract. At 20 epochs, QLSTM continues to improve, reaching 93.8% accuracy, and by 50 epochs it achieves 94.7% accuracy, 95.1% precision, and 94.7% F1-score—surpassing both classical models across all metrics. While LSTM shows a steady but limited improvement, it peaks at 88.2% accuracy before dropping at 50 epochs, suggesting overfitting. CNN shows a similar trend, slightly improving until 20 epochs but then decreasing in performance, indicating limited ability to generalize across different attack scenarios. Figure 3 (a) illustrates how the accuracy evolves as the number of epochs increases for QLSTM, LSTM, and CNN. Unlike the classical models, which tend to plateau or degrade with more training, QLSTM continues to improve, reflecting its deeper learning capacity. Figures 3 (b, c, and d) further confirm QLSTM’s superior performance, showing that it consistently achieves the highest precision, recall, and F1-score. While CNN initially maintains balanced metrics, it begins to overfit, particularly at higher epochs. LSTM retains moderate balance but lacks the expressiveness to fully distinguish between quantum attack types and normal traffic in QKD processes.
Figure 3: Performance comparison of different models: a) Accuracy vs. Epochs for Hybrid QLSTM, LSTM, and CNN models, b) Precision vs. Epochs for Hybrid QLSTM and Classical LSTM, c) Recall vs. Epochs for Hybrid QLSTM and Classical LSTM, d) F1-score vs. Epochs for Hybrid QLSTM and Classical LSTM, and e) Loss vs. Epochs for Hybrid QLSTM, LSTM, CNN models.
Figure 4: Confusion matrices for the evaluated models: (a) CNN, (b) LSTM, and (c) Hybrid QLSTM.
To further analyze model behavior, Figure 4 presents the confusion matrices for CNN, LSTM, and QLSTM at 50 epochs. QLSTM demonstrates the most consistent and accurate classification across all eight classes. It successfully detects Combined Attack (class 0), Detector Blinding Attack (class 1), Intercept and Resend Attack (class 2), Normal (class 3), PNS Attack (class 4), RNG Attack (class 5), Trojan Horse Attack (class 6), and Wavelength Dependent Trojan Attack (class 7). Most misclassifications in QLSTM are limited to overlaps between Normal and PNS Attack, which share similar photon-level patterns. In comparison, LSTM and CNN exhibit broader confusion, especially with Detector Blinding and Trojan Horse attacks. CNN also struggles more with differentiating Normal and PNS Attack, showing higher false positives and negatives. While LSTM handles Detector Blinding Attack slightly better, it still misclassifies Intercept and Resend and Normal traffic due to their overlapping behaviors. These results confirm that QLSTM is more capable of learning the complex quantum patterns needed for accurate classification in QKD environments.
Table VIII presents the final evaluation metrics after 50 epochs of training. The QLSTM achieves better performance across all considered metrics, confirming its enhanced ability to detect and classify attacks in the QKD system.
TABLE VIII: Performance Metrics Comparison across Different Models at Different Epochs
Model Epochs Accuracy Precision Recall F1-Score Test Loss
LSTM 10 85.8% 89.5% 85.8% 84.9% 0.258
20 88.2% 90.3% 88.2% 87.8% 0.239
50 85.1% 87.1% 85.5% 84.5% 0.310
CNN 10 86.8% 86.9% 86.8% 86.5% 0.236
20 86.5% 86.8% 86.5% 86.2% 0.224
50 84.9% 85.7% 84.9% 84.8% 0.303
QLSTM 10 88.3% 90.0% 88.3% 87.0% 0.216
20 93.8% 94.2% 93.8% 93.8% 0.208
50 94.7% 95.1% 94.7% 94.7% 0.189
VI Conclusion
In this work, we addressed security vulnerabilities in QKD by developing a hybrid intrusion detection system based on QLSTM. While QKD provides theoretically unbreakable encryption, its real-world implementations remain vulnerable to various attacks, such as Intercept-Resend, PNS, and Trojan Horse attacks. To enhance the security of QKD networks, this work integrates quantum-enhanced learning into a deep learning-based intrusion detection framework.
A key contribution of this work is the development of a realistic QKD dataset due to the lack of publicly available datasets for QKD attack detection. The dataset includes essential quantum security metrics, such as QBER, measurement entropy, signal and decoy loss rates, and time-based metrics, ensuring an accurate representation of practical QKD conditions. The dataset incorporates three attack scenarios: intercept-and-resend, PNS, Trojan Horse, wavelength-dependent Trojan Horse, RNG, Detector Blinding, and combined attacks.
This work introduces a QLSTM model for detecting attacks on QKD systems by capturing the temporal patterns in quantum metrics. The model’s performance was evaluated against classical deep learning models, including LSTM, and CNN. The results showed that QLSTM not only achieved high accuracy of 93.7% after 50 training epochs on the proposed dataset, but also demonstrated stronger generalization capabilities. Overall, the results demonstrate the effectiveness of the proposed QLSTM model for intrusion detection in QKD systems. This promising performance suggests that hybrid quantum-classical approaches could play an important role in enhancing the security of future QKD networks.
Despite the promising results, the proposed model has certain limitations, which provide several areas for future investigations. First, the current training process utilized fixed hyperparameters, leaving room for further exploration through fine-tuning. Adjusting the learning rate, batch size, and optimizer settings could enhance the model’s convergence and generalization. Second, the current evaluation focused on normal QKD and three attack scenarios. Considering more attack scenarios within the dataset would enable a more comprehensive evaluation of the model’s robustness. Third, the current dataset simulates a 20-bit key, so future work should aim to generate datasets for 256-bit keys to better reflect real-world cryptographic requirements. Fourth, adopting alternative or more advanced models, such as Gated Recurrent Units (GRUs), or quantum-enhanced GRUs, could offer improvements in both performance and computational efficiency. Lastly, benchmarking against proposed models from the literature would provide further insights into the strengths and weaknesses of the QLSTM model in real-world applications.